# multi-sr.bib

### v1

@@ 1 Introduction
Super-Resolution Problem. The problem of super-resolution (SR) is to recover the fine details of a signal from inherently low resolution measurements in the frequency domain. The SR problem has a variety of applications in signal processing, imaging, optics, inverse scattering, and data analysis problems [Bertero1996, lindberg2012]. In the last years, a particular SR model has received considerable attention, where the signal is modelled as a sparse measure [batenkov2021a]. Let $\mathbb{T}^{d}:=[-\pi,\pi)^{d}\equiv\left(\mathbb{R}\mod 2\pi\right)^{d}$, let $\delta(\cdot)$ denote the Dirac delta distribution, and consider the signal
$f(t):=\sum_{j=1}^{n}\alpha_{j}\delta(t-x_{j}),\quad t,x_{j}\in\mathbb{T}^{d},\quad\alpha_{j}\in\mathbb{C},$ (1)
which we aim to recover. We refer to $\{x_{j}\}$ as the ”nodes” and to $\{\alpha_{j}\}$ as the ”coefficients” or “amplitudes”. The noisy measurements are given by
@@
One primary objective in theory of super-resolution is to define optimal bounds for reconstruction errors, commonly referred to as min-max error bounds, or the computational resolution limits [donoho1992a, liu2021, Li&Liao2020, demanet2015, batenkov2021a]. These bounds are achieved by the (non-tractable) most effective algorithm under the worst-case scenario. Furthermore, these stability bounds serve to validate the optimality of different tractable solution methods, offering assurances regarding the methods’ performance [li2020a, katz2023b, katz2024b]. It has been showed previously that the min-max bounds are related to the smallest singular value of the Vandermonde matrix $U$ of the system [demanet2015], which is defined as
@@
$U(x_{1},\dots,x_{n};{\cal G}_{N}):=\big{[}\exp{\imath\langle\omega,x_{j}\rangle}\big{]}_{\omega\in{\cal G}_{N}}^{j=1,\dots,n}\in\mathbb{C}^{|{\cal G}_{N}|\times n}.$ (3)
@@
Extensive research has been conducted on the one-dimensional scenario, covering both well-separated and super-resolution regimes [batenkov2021, Li&Liao2020, cuyt2018, petz2021]. Let $\ell$ be the number of nodes in the largest cluster. Then, in the super-resolution regime, the smallest singular value of the Vandermonde matrix scales like $\text{SRF}^{1-\ell}$, resulting in “on-grid” min-max bounds to be of order $\text{SRF}^{2\ell-1}\epsilon$ [batenkov2018, Li&Liao2020], (cf. similar off-grid bounds [batenkov2021a, liu2022]). In the multidimensional context, several results are available for the well-separated regime as well as the super-resolution regime [liu2021, kunis2021, kunis2020a, garcia2020, poon2019], under specific conditions related to the unknown nodes. A key distinction between the one-dimensional and multidimensional cases lies in the spectrum of the Vandermonde matrix, wherein the geometry of the nodes plays a crucial role. As we shall show in this paper, and consistent with the observations in [kunis2020a], it is not solely the distance between the nodes that determines the spectral properties of the Vandermonde matrix, at least in the near-colliding limit, but also the algebraic variety on which these nodes are situated.
@@
Kernel matrices in the flat limit. Following the previous discussion, let us consider the Gramian matrix $D_{N}:=\frac{1}{{(2N)}^{d}}U^{*}U$, which is the “kernel matrix” for the multidimensional Dirichlet kernel. Kernel matrices are of importance in various fields, including scattered data approximation and machine learning, see e.g. [usevich2021, wathen2015spectral] and references therein. In the context of super-resolution, if all the nodes $\mathcal{X}=\{x_{j}\}_{j=1}^{n}$ form a single cluster and $\text{SRF}\gg 1$, then the kernel matrix $D_{N}$ can be considered in the so-called “flat limit”, a term introduced in [driscoll2002] in the context of radial basis function (RBF) interpolation. The flat limit was investigated recently in various publications, from among those [usevich2021] and [wathen2015spectral] being most relevant to our work. While [wathen2015spectral] deals exclusively with RBF kernels and derives asymptotic eigenvalue scaling, [usevich2021] extends those results to arbitrary smooth kernels, albeit under the assumption of the points not lying on any low-dimensional algebraic variety. The approach from [wathen2015spectral] is primarily based on Micchelli’s lemma [micchelli1986], and we have adopted this approach previously in the one-dimensional case [batenkov2021].
@@
Contributions. The main result of this paper, Theorem 1, derives the asymptotic scaling of eigenvalues of a kernel matrix induced by an infinitely smooth kernel in the flat limit, generalizing the corresponding result of [usevich2021] to arbitrary geometry. As a corollary, we derive the asymptotic scaling of the eigenvalues of $D_{N}$, leading to the appropriate scaling of the singular values of the Vandermonde matrix (3), when the nodes are arranged within a single cluster. In this context, a ”single cluster” is characterized by the condition $N\Delta\ll 1$. We therefore obtain a generalization of the results from [kunis2021] which considered mainly the “single line” geometry, as well as the extension of [batenkov2021, Theorem 2.3] to the multivariate case. See also Remark 6.
@@
In more detail, suppose that $x_{j}=\Delta y_{j}$ with $\epsilon=N\Delta$, and take $\epsilon\to 0$. We show that the number of eigenvalues of $D_{N}(\{x_{1},\dots,x_{n}\})$ decaying like $\epsilon^{2k}$, is precisely equal to a certain number $t_{k}$ which depends on the algebraic properties of the original nodes $\{y_{j}\}$ (see Theorem 5). In particular, when all the nodes belong to a one-dimensional affine subspace (a “single-line” geometry), then $t_{k}=1$, which implies $\sigma_{\min}(U)\approx N^{\frac{d}{2}}\epsilon^{n-1}$ (as derived in [kunis2021]). On the other hand, when $\{y_{j}\}$ are in a general position, then, as long as $n\geq{k+d\choose d}$, we have $t_{k}={k+d-1\choose d-1}$, where $d$ is the dimension of the space. These numbers coincide for $d=1$, but may differ drastically for $d>1$. For example, if $d=2$ then $\sigma_{\min}(U)$ may be as large as $N\epsilon^{\mathcal{O}(\sqrt{n})}$. Our results are slightly more general and provide the eigenvalue scaling for arbitrary (symmetric) sampling sets $\omega\in\mathcal{S}$, with the upper bounds the same as described above; however we can prove matching lower bounds only for certain $\mathcal{S}$ satisfying the so-called Geometric Characterization Condition (in particular satisfied by ${\cal G}_{N}$), see Definition 8 below.
@@
To demonstrate the relevance of our findings to the multivariate SR problem, in Section 6 we also report on a numerical study of reconstructing the signals from noisy measurements using the nonlinear least squares (NLS) method and the multidimensional ESPRIT method [sahnoun2017]. Our results indeed suggest that the asymptotic conditioning of the problem may vary from the “worst-case” single-line scenario (previously considered in [liu2021]), to the “best-case” general position scenario. We are therefore confident that our findings will enhance the understanding of super-resolution stability in high-dimensional settings and can contribute to the analysis of current multivariate SR recovery methods such as [diederichs2022, pena2024, sauer2017, sauer2018, sahnoun2017]. Furthermore, we believe these results can serve as an important step towards establishing min-max error bounds in the general multidimensional scenario, when the nodes form multiple clusters (where some kind of multidimensional “confluent” Vandermonde matrices may be required, cf. [batenkov2013a, batenkov2013c, batenkov2018-genspikes, batenkov2023]).
@@
Organization of the paper. In Section 2 we establish some notation and definitions. In Section 3 we present our main result, Theorem 1, which describes the asymptotic scaling of the eigenvalues of kernel matrices in the flat limit. In Section 4 we specialize the results to the Dirichlet kernel matrix, showing, in particular, that the bounds of Theorem 1 are tight. In Section 5 we derive the corresponding scaling of the singular values of the Vandermonde matrix, which requires additional technical arguments. Finally, numerical results are presented in Section 6.
@@ 2 Preliminaries
We adapt some definitions and notation from [usevich2021]. For a multi-index $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{Z}^{d}_{+}$, denote
@@
For $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\boldsymbol{\alpha}\in\mathbb{Z}^{d}_{+}$, the monomial $\boldsymbol{x}^{\boldsymbol{\alpha}}$ is defined as $\boldsymbol{x}^{\boldsymbol{\alpha}}:=x_{1}^{\alpha_{1}}\cdot\cdot\cdot x_{d}^{\alpha_{d}}$, and its (total) degree is $|\boldsymbol{\alpha}|$. Let $\prec$ be a fixed graded order on the set of multi-indices +$\mathbb{Z}^{d}_{+}$ (or, equivalently, on the set of monomials), for example a graded lexicographic order or a graded reverse lexicographic order. It will always be implied in the following.
Definition 1.
@@
Definition 2.
For a kernel $\mathcal{K}\in\mathcal{C}^{(r,r)}(\Omega)$, a finite set ${\cal X}\subset\Omega$ and $\epsilon>0$ we define the scaled kernel $\mathcal{K}_{\epsilon}(\boldsymbol{x},\boldsymbol{y}):=\mathcal{K}(\epsilon\boldsymbol{x},\epsilon\boldsymbol{y})$ and the kernel matrix
$K_{\epsilon}({\cal X}):=\big{[}\mathcal{K}_{\epsilon}(\boldsymbol{x},\boldsymbol{x^{\prime}})\big{]}_{\boldsymbol{x},\boldsymbol{x^{\prime}}\in{\cal X}}\in\mathbb{R}^{|{\cal X}|\times|{\cal X}|}.$ (5)
Definition 3 (Multivariate Vandermonde matrix).
@@
Notice that $V$ is a real Vandermonde matrix, as opposed to the (transpose of the) complex Vandermonde matrix $U$ defined in (3). Let us further denote $V_{\leq k}:=V_{\mathbb{P}_{k}}({\cal X})$ and $V_{k}:=V_{\mathbb{H}_{k}}({\cal X})$. The matrix $V_{\leq k}\in\mathbb{R}^{n\times p_{k}}$ can be partitioned into $k+1$ matrices $V_{k}\in\mathbb{R}^{n\times h_{k}}$ arranged by increasing degree:
@@
Definition 5 (Discrete moment order, [schabak2005]).
@@
$\ker(V_{\leq m}({\cal X})^{T})=\emptyset\iff\operatorname{rank}(V_{\leq m}({\cal X}))=n.$
Note that such $m$ exists (is finite). Indeed, if ${\cal X}=\{x_{1},\dots,x_{n}\}$, then $\mu({\cal X})\leq n-1$ because the following Lagrange interpolation polynomials, each of total degree $n-1$, are linearly independent:
@@
where $H_{j}$ is some affine hyperplane containing $x_{j}$ and not containing every other $x_{k},k\neq j$. Cf. [schabak2005, Definition 3],[GascaMariano2000] and [wathen2015spectral].
@@ 3 Eigenvalues of smooth kernel matrices in the flat limit
Theorem 1.
Let $m=\mu({\cal X})$ be the discrete moment order of the set ${\cal X}$. For any infinitely smooth kernel $\mathcal{K}\in\mathcal{C^{(\infty,\infty)}}(\Omega)$ with $\mathbf{0}\in\Omega$, and small enough $\epsilon$, the eigenvalues of $K_{\epsilon}({\cal X})$ split into $m+1$ groups
@@
Remark 1.
Theorem (1) is valid also for kernels in $\mathcal{C}^{(m+1,m+1)}$.
@@
We start by writing the Taylor expansion of $K_{\epsilon}$ around 0 as equation (55) in [usevich2021]:
@@
where $W_{i}(\epsilon)=O(1)$, $W=W^{\mathcal{K}}_{\leq m}$ and $\Delta_{m}(\epsilon)=\operatorname{diag}(1,\epsilon I_{h_{1}},\dots,\epsilon^{m}I_{h_{m}})\in\mathbb{R}^{p_{m}\times p_{m}}$. For the full derivation of (9), see section A in the Appendix.
Let $V=V_{\leq m}$ where $V=[V_{0}\dots V_{m}]=[V_{\leq m-1}\quad V_{m}]$, and $t_{k}:=\operatorname{rank}(V_{\leq k})-\operatorname{rank}(V_{\leq k-1})$ as defined in the theorem. In [usevich2021] only the generic case $t_{k}=h_{k}-h_{k-1}$ is considered. Here we generalize Theorem 6.1 in [usevich2021] to the case where the multivariate Vandermonde matrices $V_{\leq k}$ are not necessarily full rank.
Consider the full QR decomposition of $V_{\leq m-1}$ with column pivoting, as elaborated in Section 5.4.2, specifically, formula (5.4.6) in the book [golub2013]:
@@
We have $[Q_{\leq m-1}\quad Q_{\perp}]^{T}V_{\leq m}P=\begin{bmatrix}R&Q_{\leq m-1}^{T}V_{m}\\ 0&Q_{\perp}^{T}V_{m}\end{bmatrix}$ and $Q_{\perp}^{T}V_{\leq m}P=[0\quad Q_{\perp}^{T}V_{m}]$ where $P:=\begin{bmatrix}P_{m-1}&0\\ 0&I_{h_{m}}\end{bmatrix}$. Note that for every permutation matrix, we have $P^{-1}=P^{T}$.
@@
where $\tilde{V}_{\leq m}:=V_{\leq m}P$ , $\tilde{\Delta}_{m}:=P^{T}\Delta_{m}P=\operatorname{diag}(\epsilon^{i_{1}}I_{h_{i_{1}}},...,\epsilon^{i_{m}}I_{h_{i_{m}}},\epsilon^{m}I_{h_{m}})$ due to Lemma 7, $\tilde{W}:=P^{T}WP$, $\tilde{W}_{1}(\epsilon):=P^{T}W_{1}(\epsilon)$ and $\tilde{W}_{2}(\epsilon):=W_{2}(\epsilon)P$.
@@
$Q_{\perp}^{T}K_{\epsilon}Q_{\perp}$ $=\quad[0\quad Q_{\perp}^{T}V_{m}]\tilde{\Delta}\tilde{W}\tilde{\Delta}[0\quad Q_{\perp}^{T}V_{m}]^{T}$
$\quad+\epsilon^{m+1}[0\quad Q_{\perp}^{T}V_{m}]\tilde{\Delta}\tilde{W}_{1}Q_{\perp}$
$\quad+\epsilon^{m+1}Q_{\perp}^{T}\tilde{W}_{2}\tilde{\Delta}[0\quad Q_{\perp}^{T}V_{m}]^{T}$
@@
For any $k=1,...,m-1$, let $V_{\leq k-1}P_{k-1}=[Q_{\leq k-1}\quad Q_{\perp,k-1}]\begin{bmatrix}R_{\leq k-1}\\ 0\end{bmatrix}$ be the full QR decomposition of $V_{\leq k}$ with pivoting, where $Q_{\leq k-1}\in\mathbb{R}^{n\times[n-(t_{k}+\dots+t_{m})]}$, $Q_{\perp,k-1}\in\mathbb{R}^{n\times[t_{k}+\dots+t_{m}]}$ and $R_{\leq k-1}\in\mathbb{R}^{[n-(t_{k}+\dots+t_{m})]\times p_{k-1}}$.

### v2

# Spectral Properties of Infinitely Smooth Kernel Matrices in the Single Cluster Limit, with Applications to Multivariate Super-Resolution

@@ Abstract
Keywords— kernel matrices, flat limit, dirichlet kernel, eigenvalues, multivariate Vandermonde matrices, super-resolution
MSC— 15A18, 65T40, 65F20, 47A55, 47A75, 47B34
@@ 1 Introduction
Super-Resolution Problem. The problem of super-resolution (SR) is to recover the fine details of a signal from inherently low resolution measurements in the frequency domain. The SR problem has a variety of applications in signal processing, imaging, optics, inverse scattering, and data analysis problems [undefh, undefaa]. In the last years, a particular SR model has received considerable attention, where the signal is modelled as a sparse measure [undefe]. Let $\mathbb{T}^{d}:=[-\pi,\pi)^{d}\equiv\left(\mathbb{R}\mod 2\pi\right)^{d}$, let $\delta(\cdot)$ denote the Dirac delta distribution (in the sense of generalized functions [undefao]), and consider the signal
$f(t):=\sum_{j=1}^{n}\alpha_{j}\delta(t-x_{j}),\quad t,x_{j}\in\mathbb{T}^{d},\quad\alpha_{j}\in\mathbb{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$ (1)
We refer to $\{x_{j}\}$ as the ”nodes” and to $\{\alpha_{j}\}$ as the ”coefficients” or “amplitudes”. The noisy measurements are given by
@@
The problem of super-resolution is, given discrete noisy Fourier measurements (2), recover the parameters of model (1), i.e. $\{\alpha_{j},x_{j}\}$.
@@
One primary objective in theory of super-resolution is to define optimal bounds for reconstruction errors, commonly referred to as min-max error bounds, or the computational resolution limits [undefm, undefab, undefy, undefk, undefe]. These bounds are achieved by the (intractable) “oracle” algorithm under the worst-case scenario. Furthermore, these stability bounds serve to validate the optimality of different tractable solution methods, offering assurances regarding the methods’ performance [undefz, undeft, undefu]. It has been shown previously that the min-max bounds are related to the smallest singular value of the Vandermonde matrix $U$ of the system [undefk], which is defined as
@@
$U(x_{1},\dots,x_{n};{\cal G}_{N}):=\big[\exp{\imath\langle\omega,x_{j}\rangle}\big]_{\omega\in{\cal G}_{N}}^{j=1,\dots,n}\in\mathbb{C}^{|{\cal G}_{N}|\times n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\ \langle\omega,x_{j}\rangle:=\omega^{*}x_{j}=\sum_{k=1}^{d}\omega_{k}^{*}x_{j,k},}$ (3)
where $\omega^{*}$ denotes the conjugate transpose of $\omega$.
@@
Extensive research has been conducted on the one-dimensional $(d=1)$ scenario, covering both well-separated and super-resolution regimes [undefg, undefy, undefj, undefag]. Let $\ell$ be the number of nodes in the largest cluster. Then, in the super-resolution regime, the smallest singular value of the Vandermonde matrix scales like $\text{SRF}^{1-\ell}$ whenever $N\Delta\geq c(n)$, resulting in “on-grid” min-max bounds to be of order $\text{SRF}^{2\ell-1}\epsilon$ [undeff, undefy], (cf. similar off-grid bounds [undefe, undefac]). In the multidimensional context, several results are available for the well-separated regime as well as the super-resolution regime [undefab, undefw, undefv, undefp, undefah], under specific conditions related to the unknown nodes. A key distinction between the one-dimensional and multidimensional cases lies in the spectrum of the Vandermonde matrix, wherein the geometry of the nodes plays a crucial role. As we shall show in this paper, and consistent with the observations in [undefv], it is not solely the distance between the nodes that determines the spectral properties of the Vandermonde matrix, at least in the near-colliding limit, but also the algebraic variety on which these nodes are situated.
@@
Kernel matrices in the flat limit. Following the previous discussion, let us consider the Gramian matrix $D_{N}:=\frac{1}{{(2N)}^{d}}U^{*}U$, which is the “kernel matrix” for the multidimensional Dirichlet kernel. Kernel matrices are of importance in various fields, including scattered data approximation and machine learning, see e.g. [undef, undefaq, undefap] and references therein. In the context of super-resolution, if all the nodes $\mathcal{X}=\{x_{j}\}_{j=1}^{n}$ form a single cluster and $\text{SRF}\gg 1$, then the kernel matrix $D_{N}$ can be considered in the so-called “flat limit”, a term introduced in [undefn] in the context of radial basis function (RBF) interpolation. The flat limit was investigated recently in various publications, from among those [undef] and [undefaq] being most relevant to our work. While [undefaq] deals exclusively with RBF kernels and derives asymptotic eigenvalue scaling, [undef] extends those results to arbitrary smooth kernels, albeit under the assumption of the points not lying on any low-dimensional algebraic variety. The approach from [undefaq] is primarily based on Micchelli’s lemma [undefad], and we have adopted this approach previously in the one-dimensional case [undefg].
@@
Contributions. The main result of this paper, Theorem 3.1, derives the asymptotic scaling of eigenvalues of a kernel matrix induced by an infinitely smooth kernel in the flat limit. This generalizes the corresponding result of [undef] to arbitrary geometry, and the result in Theorem 8 of [undefaq] to arbitrary smooth kernels.. As a corollary, we derive the asymptotic scaling of the eigenvalues of $D_{N}$, leading to the appropriate scaling of the singular values of the Vandermonde matrix (3), when the nodes are arranged within a single cluster. In this context, a ”single cluster” is characterized by the condition $N\Delta\ll 1$. We therefore obtain a generalization of the results from [undefw] which considered mainly the “single line” geometry, as well as the extension of [undefg, Theorem 2.3] to the multivariate case. See also Remark 5.14.
@@
In more detail, suppose that $x_{j}=\Delta y_{j}$ with $\epsilon=N\Delta$, and take $\epsilon\to 0$. We show that the number of eigenvalues of $D_{N}(\{x_{1},\dots,x_{n}\})$ decaying like $\epsilon^{2k}$, is precisely equal to a certain number $t_{k}$ which depends on the algebraic properties of the original nodes $\{y_{j}\}$ (see Theorem 5.3). In particular, when all the nodes belong to a one-dimensional affine subspace (a “single-line” geometry), then $t_{k}=1$, which implies $\sigma_{\min}(U)\approx N^{\frac{d}{2}}\epsilon^{n-1}$ (as derived in [undefw]). On the other hand, when $\{y_{j}\}$ are in a general position, then, as long as $n\geq{k+d\choose d}$, we have $t_{k}={k+d-1\choose d-1}$, where $d$ is the dimension of the space. These numbers coincide for $d=1$, but may differ drastically for $d>1$. For example, if $d=2$ then $\sigma_{\min}(U)$ may be as large as $N\epsilon^{\mathcal{O}(\sqrt{n})}$. Our results are slightly more general and provide the eigenvalue scaling for arbitrary (symmetric) sampling sets $\omega\in\mathcal{S}$, with the upper bounds the same as described above; however we can prove matching lower bounds only for certain $\mathcal{S}$ satisfying that either satisfy the the so-called Geometric Characterization Condition (in particular satisfied by ${\cal G}_{N}$), see Definition 4.8 below or are ’lower sets’ as described in Definition 4.16.
@@
To demonstrate the relevance of our findings to the multivariate SR problem, in Section 6 we also report on a numerical study of reconstructing the signals from noisy measurements using the nonlinear least squares (NLS) method and the multidimensional ESPRIT method [undefaj]. Our results indeed suggest that the asymptotic conditioning of the problem may vary from the “worst-case” single-line scenario (previously considered in [undefab]), to the “best-case” general position scenario. We are therefore confident that our findings will enhance the understanding of super-resolution stability in high-dimensional settings and can contribute to the analysis of current multivariate SR recovery methods such as [undefl, undefaf, undefal, undefam, undefaj]. Furthermore, we believe these results can serve as an important step towards establishing min-max error bounds in the general multidimensional scenario, when the nodes form multiple clusters (where some kind of multidimensional “confluent” Vandermonde matrices may be required, cf. [undefb, undefa, undefc, undefd]).
@@
Organization of the paper. In Section 2 we establish some notation and definitions. In Section 3 we present our main result, Theorem 3.1, which describes the asymptotic scaling of the eigenvalues of kernel matrices in the flat limit. In Section 4 we specialize the results to the Dirichlet kernel matrix, showing, in particular, that the bounds of Theorem 3.1 are tight. In Section 5 we derive the corresponding scaling of the singular values of the Vandermonde matrix, which requires additional technical arguments. Finally, numerical results are presented in Section 6.
@@ 2 Preliminaries
We adapt some definitions and notation from [undef]. For a multi-index $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{Z}^{d}_{+}$, denote
@@
For $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\boldsymbol{\alpha}\in\mathbb{Z}^{d}_{+}$, the monomial $\boldsymbol{x}^{\boldsymbol{\alpha}}$ is defined as $\boldsymbol{x}^{\boldsymbol{\alpha}}:=x_{1}^{\alpha_{1}}\cdot\cdot\cdot x_{d}^{\alpha_{d}}$, and its (total) degree is $|\boldsymbol{\alpha}|$. Let $\prec$ be a fixed graded order on the set of multi-indices $\mathbb{Z}^{d}_{+}$ (or, equivalently, on the set of monomials), for example a graded lexicographic order or a graded reverse lexicographic order. The ordering $\prec$ is used throughout the sequel.
Definition 2.1.
@@
Definition 2.2.
For a kernel $\mathcal{K}\in\mathcal{C}^{(r,r)}(\Omega)$, a finite set ${\cal X}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}\}}\subset\Omega$ and $\epsilon>0$ we define the scaled kernel $\mathcal{K}_{\epsilon}(\boldsymbol{x},\boldsymbol{y}):=\mathcal{K}(\epsilon\boldsymbol{x},\epsilon\boldsymbol{y})$ and the kernel matrix
$K_{\epsilon}({\cal X}):=\big[\mathcal{K}_{\epsilon}(\boldsymbol{x},\boldsymbol{x^{\prime}})\big]_{\boldsymbol{x},\boldsymbol{x^{\prime}}\in{\cal X}}\in\mathbb{R}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n\times n}.$ (5)
Definition 2.3 (Multivariate Vandermonde matrix).
@@
Note that $V$ is a multivariate real Vandermonde matrix with monomial entries, in contrast to the complex Vandermonde matrix $U$ defined in (3), whose entries are of the form $e^{\imath\langle\omega,x\rangle}$. Let us further denote $V_{\leq k}:=V_{\mathbb{P}_{k}}({\cal X})$ and $V_{k}:=V_{\mathbb{H}_{k}}({\cal X})$. The matrix $V_{\leq k}\in\mathbb{R}^{n\times p_{k}}$ can be partitioned into $k+1$ matrices $V_{k}\in\mathbb{R}^{n\times h_{k}}$ arranged by increasing degree:
@@
Definition 2.5 (Discrete moment order, [undefan]).
@@
$\ker(V_{\leq m}({\cal X})^{T})=\emptyset$
holds, or equivalently,
$\operatorname{rank}(V_{\leq m}({\cal X}))=n$
holds.
Note that such $m$ exists (is finite). Indeed, if ${\cal X}=\{x_{1},\dots,x_{n}\}$, then $\mu({\cal X})\leq n-1$ because the following Lagrange interpolation polynomials, each of total degree n1n-1, are linearly independent:
@@
where $H_{j}$ is some affine hyperplane containing $x_{j}$ and not containing every other $x_{k},k\neq j$. Cf. [undefan, Definition 3],[undefq] and [undefaq].
@@ 3 Eigenvalues of smooth kernel matrices in the flat limit
Theorem 3.1.
[…]
@@
Remark 3.2.
Theorem (3.1) is valid also for kernels in $\mathcal{C}^{(m+1,m+1)}$.
@@
We start by writing the Taylor expansion of $K_{\epsilon}$ around 0 as equation (55) in [undef]:
@@
where $_{i}(\epsilon)=O(1), W=W$$W^{\mathcal{K}}_{\leq m}$ and $\Delta_{m}(\epsilon)=\operatorname{diag}(1,\epsilon I_{h_{1}},\dots,\epsilon^{m}I_{h_{m}})\in\mathbb{R}^{p_{m}\times p_{m}}$. For the full derivation of (9), see section A in the Appendix.
Let $V=V_{\leq m}$ where $V=[V_{0}\dots V_{m}]=[V_{\leq m-1}\quad V_{m}]$, and $t_{k}:=\operatorname{rank}(V_{\leq k})-\operatorname{rank}(V_{\leq k-1})$ as defined in the theorem. In [undef] only the generic case $t_{k}=h_{k}-h_{k-1}$ is considered. Here we generalize Theorem 6.1 in [undef] to the case where the multivariate Vandermonde matrices $V_{\leq k}$ are not necessarily full rank.
Consider the full QR decomposition of $V_{\leq m-1}$ with column pivoting, as elaborated in Section 5.4.2, specifically, formula (5.4.6) in the book [undefr]:
@@
We have $Q_{\leq m-1}\quad Q_{\perp}]^{T}V_{\leq m}P=\begin{bmatrix}R&Q_{\leq m-1}^{T}V_{m}\\ 0&Q_{\perp}^{T}V_{m}\end{bmatrix}$ and $Q_{\perp}^{T}V_{\leq m}P=[0\quad Q_{\perp}^{T}V_{m}]$ where $P:=\begin{bmatrix}P_{m-1}&0\\ 0&I_{h_{m}}\end{bmatrix}$. Note that for every permutation matrix, we have $P^{-1}=P^{T}$.
@@
where $\tilde{V}_{\leq m}:=V_{\leq m}P$ , $\tilde{\Delta}_{m}:=P^{T}\Delta_{m}P={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{diag}(\epsilon^{i_{1}},...,\epsilon^{i_{q_{m-1}}},\epsilon^{m}I_{h_{m}})}$, $q_{m-1}:=\sum_{i=0}^{m-1}h_{i}$ due to Lemma B.1, $\tilde{W}:=P^{T}WP$, $\tilde{W}_{1}(\epsilon):=P^{T}W_{1}(\epsilon)$ and $\tilde{W}_{2}(\epsilon):=W_{2}(\epsilon)P$.
@@
$Q_{\perp}^{T}K_{\epsilon}Q_{\perp}$ $=\quad[0\quad Q_{\perp}^{T}V_{m}]\tilde{\Delta}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}\tilde{W}\tilde{\Delta}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}[0\quad Q_{\perp}^{T}V_{m}]^{T}$
$\quad+\epsilon^{m+1}[0\quad Q_{\perp}^{T}V_{m}]\tilde{\Delta}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}\tilde{W}_{1}Q_{\perp}$
$\quad+\epsilon^{m+1}Q_{\perp}^{T}\tilde{W}_{2}\tilde{\Delta}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}[0\quad Q_{\perp}^{T}V_{m}]^{T}$
@@
For any k=1,,m1k=1,...,m-1, let $V_{\leq k-1}P_{k-1}=[Q_{\leq k-1}\quad Q_{\perp,k-1}]\begin{bmatrix}R_{\leq k-1}\\ 0\end{bmatrix}$ be the full QR decomposition of $V_{\leq k}$ with pivoting, where $Q_{\leq k-1}\in\mathbb{R}^{n\times[n-(t_{k}+\dots+t_{m})]}$, $Q_{\perp,k-1}\in\mathbb{R}^{n\times[t_{k}+\dots+t_{m}]}$ and $R_{\leq k-1}\in\mathbb{R}^{[n-(t_{k}+\dots+t_{m})]\times p_{k-1}}$.